OpenStudy (anonymous):

Bob rolls a fair six-sided die each morning. If Bob rolls a composite number, he eats sweetened cereal. If he rolls a prime number, he eats unsweetened cereal. If he rolls a 1, then he rolls again. In a non leap year, what is the expected value of the difference between the number of days Bob eats unsweetened cereal and the number of days he eats sweetened cereal?

OpenStudy (anonymous):

what about 1?

OpenStudy (anonymous):

oh he rolls again

OpenStudy (anonymous):

primes are 2, 3, 5, composite 4, 6, so he gets prime \(\frac{3}{5}\) of the time, composite \(\frac{2}{5}\) of the time

OpenStudy (anonymous):

\[\frac{3}{5}\times 365-\frac{2}{5}\times 365\] will give you the expected value of the difference

OpenStudy (anonymous):

Thanks